1) the bond dissociation enthalpies for H2 and Hcl are 435kj/mol and 432 kj/mol respectively if the Hr^o for the formation of HCl (g) from H2 and Cl2 is -184.6 kj/mol what is the bond dissociation enthalpies for C2?


2) A 20.9g of mercury (Ii) oxide (HgO , MW= 216.6g/mol) is heated to form the decomposition product Hg and O2 gas. what volume of O2 is produced?

answer all please with work. thank u
Answer #1

Solution 0 H-H + cl-ce 2 H -(l SHaxn = -184-6 kJ/mol Type Bonds broken Number 1 - 1 H-H Kilmal 435 = 435kJ/mol x =XkJ/mol cl-moles of HgO = mass of Hgo Molar mass of Hgo 20.99 216.691mal = 0.0965 mal From the above chemical reaction it is clear 2 mal